Opened 6 years ago

Last modified 5 years ago

#455 assigned flight processing

EUFAR12/02, flight day 240b/2012, North Iceland

Reported by: knpa Owned by: knpa
Priority: alpha 4 high Milestone: 2012 data processing completion
Component: Archiving Keywords:
Cc: Other processors: anhi tipo edfi

Description (last modified by edfi)

Data location: /users/rsg/arsf/arsf_data/2012/flight_data/greenland_iceland/EUFAR12_02-2012_240b_North_Iceland

Data arrived from ARSF via SATA disk on 24/09/2012

EUFAR12/02 Scientific objective: unknown

EUFAR12/02 Priority: a4h

PI: Susan Conway

Sensors:

Camera (20/12/2012)
Eagle (20/12/2012)
Hawk (20/12/2012)
LiDAR (20/12/2012)

Change History (25)

comment:1 Changed 6 years ago by knpa

  • Description modified (diff)
  • Priority changed from alpha 4 medium to alpha 4 high

comment:2 Changed 6 years ago by knpa

  • Description modified (diff)

comment:3 Changed 6 years ago by anhi

Having problems doing the nav processing with the basestations provided by the PI

comment:4 Changed 6 years ago by anhi

Disabled GLONASS in the nav processing because it was failing.

comment:5 Changed 6 years ago by anhi

LiDAR
Pitch:-0.0032
Roll:0.00038

comment:6 Changed 6 years ago by anhi

LiDAR delivery created

comment:7 Changed 6 years ago by edfi

I think there needs to be a comment about the hole in line (I think) 17, just the distance exceeding the maximum the scanner was set up for. Everything else looks fine

comment:8 Changed 6 years ago by anhi

Hole was caused by cloud, have added a comment about it and remade the Read Me.

comment:9 Changed 6 years ago by edfi

Starting second check of LiDAR

comment:10 Changed 6 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

Beginning RCD processing.

comment:11 Changed 6 years ago by knpa

  • Other processors set to anhi

comment:12 Changed 6 years ago by knpa

There are two set of logfiles. The second lot are relevant to the images in the project (121115).

comment:13 Changed 6 years ago by tipo

Started hyperspectral processing

comment:14 Changed 6 years ago by tipo

SCT values:

LineEagleHawk
1-0.02-0.05
2-0.02-0.05
3-0.02-0.05
4-0.02-0.05
5-0.02-0.05
6-0.02-0.05
7-0.02-0.05
8-0.02-0.05
9-0.02-0.05
101.85-0.05
11-0.02-0.05
12-0.02-0.05
13-0.08-0.05
14-0.02-0.05
15-0.02-0.05
16-0.02-0.05
17-0.02-0.05
18-0.02-0.06
19-0.02-0.05
20-0.02-0.05
Last edited 6 years ago by tipo (previous) (diff)

comment:15 Changed 6 years ago by tipo

Hyperspectral processing complete.

comment:16 Changed 6 years ago by edfi

Started hyperspectral delivery check

comment:17 Changed 6 years ago by edfi

Hyperspectral delivery check complete, changed logsheet from handwritten to electronic all else fine.

comment:18 Changed 6 years ago by knpa

  • Other processors changed from anhi to anhi tipo edfi

RCD: anhi has created the tiffs

comment:19 Changed 6 years ago by knpa

RCD ready for delivery check.

comment:20 Changed 6 years ago by tipo

Camera delivery check complete. No problems found.

comment:21 Changed 6 years ago by edfi

  • Description modified (diff)

comment:22 Changed 6 years ago by edfi

Dispatched to Susan Conway 20/12/2012

comment:23 Changed 6 years ago by knpa

  • Component changed from Processing: general to Archiving

Beginning archiving.

comment:24 Changed 6 years ago by knpa

Made available to NEODC.

comment:25 Changed 5 years ago by knpa

archived by NEODC.

Note: See TracTickets for help on using tickets.